# SegAssess: Panoramic quality mapping for robust and transferable unsupervised segmentation assessment

### v1

@@ Abstract
High-quality image segmentation is fundamental to pixel-level geospatial analysis in remote sensing, necessitating robust segmentation quality assessment (SQA), particularly in unsupervised settings lacking ground truth. Although recent deep learning (DL) based unsupervised SQA methods show potential, they often suffer from coarse evaluation granularity, incomplete assessments, and poor transferability. To overcome these limitations, this paper introduces Panoramic Quality Mapping (PQM) as a new paradigm for comprehensive, pixel-wise SQA, and presents SegAssess, a novel deep learning framework realizing this approach. SegAssess distinctively formulates SQA as a fine-grained, four-class panoramic segmentation task, classifying pixels within a segmentation mask under evaluation into true positive (TP), false positive (FP), true negative (TN), and false negative (FN) categories, thereby generating a complete quality map. Leveraging an enhanced Segment Anything Model (SAM) architecture, SegAssess uniquely employs the input mask as a prompt for effective feature integration via cross-attention. Key innovations include an Edge Guided Compaction (EGC) branch with an Aggregated Semantic Filter (ASF) module to refine predictions near challenging object edges, and an Augmented Mixup Sampling (AMS) training strategy integrating multi-source masks to significantly boost cross-domain robustness and zero-shot transferability. Comprehensive experiments across 32 datasets derived from 6 sources demonstrate that SegAssess achieves state-of-the-art (SOTA) performance and exhibits remarkable zero-shot transferability to unseen masks, establishing PQM via SegAssess as a robust and transferable solution for unsupervised SQA. The code is available at https://github.com/Yangbn97/SegAssess
@@ I Introduction
The proliferation of high-resolution remote sensing imagery from diverse sensing platforms has revolutionized geospatial analysis, enabling unprecedented spatial detail for Earth observation[1, 2]. Image segmentation (IS), which partitions pixels into semantically meaningful regions, serves as the foundational technique[3, 4, 5]. By abstracting pixel-level data into object-based masks, IS aligns with human perceptual mechanisms, offering enhanced interpretability for various tasks such as land-cover /land-use mapping[6, 7, 8], resource management[9, 10], disaster monitoring[11, 12], urban planning[13, 14], etc.
Image segmentation approaches are typically categorized into two schemes: human-centric methods, such as manual annotation or interactive segmentation, and automated model-driven solutions, including empirical algorithms or deep learning (DL) based image segmentation models. Both paradigms possess inherent limitations that necessitate rigorous quality assessment. Human-centric annotation introduces quality uncertainty due to subjective biases, variations in expertise, and inter-annotator inconsistencies[15, 16] while model-driven segmentation is sensitive to methodological choices, parameter tuning, and computational environments, leading to variability[1, 17]. As the basis of pixel-level geospatial analysis, the segmentation quality remarkably impacts the reliability and robustness of the entire workflow. This critical dependency underscores the urgent demand for robust segmentation quality assessment (SQA) frameworks to enforce the performance control in real-world image interpretation applications.
@@
There have been some efforts devoted to evaluate the segmentation quality. As shown in Fig. 1, existing SQA methods broadly fall into two categories: human-centric subjective evaluation and algorithm-based objective evaluation. Human-centric subjective assessment, the traditional and often primary paradigm, involves visual inspection by domain experts[18, 19]. While theoretically capable of capturing nuanced quality aspects, this approach faces significant practical constraints. Substantial assessment uncertainty can arise from inter-operator variability due to subjective interpretation biases, differing expertise levels, and cognitive inconsistencies[1, 20, 2]. Furthermore, the labor-intensive nature of manual evaluation requires specialized training and extensive time investments, which renders it economically impractical for large-scale applications and motivates the development of automated alternatives.
[…]
Unsupervised SQA approaches[31, 32, 33, 34, 35, 36, 37, 38, 39] offer an alternative by eliminating the dependency on reference knowledge (e.g., ground truth) and instead perform goodness evaluation via the implicit pattern analysis of segmentation masks and their corresponding images. In this scope, early unsupervised SQA works[31, 32, 40] primarily focused on evaluating the intra-segment homogeneity and inter-segment heterogeneity using expert-designed indicators. These methods typically combined empirically designed indices, such as spectral variance[41, 42, 43] and texture uniformity[44, 45], to quantify segmentation quality at the image or object level. These approaches can obtain merit performance within their targeting data distribution, but often struggle with transferability to new tasks or datasets. With the emergence of deep learning, some works managed to quantify segmentation quality by establishing score prediction (SP) networks[34, 38, 39]. Theses models simultaneously process images and unchecked segmentation masks to predict a predefined quality scores in a regression manner (Fig. 1). Though demonstrating effectiveness in certain contexts, these score-based unsupervised SQA approaches still face three key challenges: 1) Subjectivity in metric design: Quality indices are formulated based on expert domain knowledge, introducing inherent subjective biases into the assessment process; 2) Coarse evaluation granularity: Assessments operate at the image or object level, lacking pixel-wise diagnostic capabilities to localize specific defects; 3) Accessibility barriers: The abstract nature of composite metrics creates often requires domain expertise for interpretation, limiting their practical usability for common people.
@@
To address aforementioned challenges of SP models, one direction explored in recent unsupervised SQA research involves pixel-level error localization (PEL) networks[35, 36]. By directly pinpointing per-pixel inaccuracies within the segmentation mask these methods attempt to obtain pixel-wise, objective and intuitionistic assessment results rather than regressing expert-customized metrics (Fig. 1). For example,[36] employed ”correct/incorrect” binary segmentation to simply identify erroneous pixels. Meanwhile, in the remote sensing filed, AQSNet[35] was designed to generate separate masks identifying false positive (FP, wrong background) and false negative (FN, missing foreground) pixels within the mask under evaluation. However, these PEL approaches, despite offering finer spatial detail than score-based methods, introduce their own set of limitations. Binary error classification lacks the necessary nuance to distinguish between different types of mistakes. Even methods like AQSNet that differentiate FP and FN errors face significant drawbacks. As illustrated by preliminary experiments(Fig 2), AQSNet provides an incomplete assessment by overlooking correct categories, i.e.true positive (TP) and true negative (TN). Furthermore, its strategy of predicting FP and FN independently can lead to spatially overlapping error regions (pink areas in Fig. 2), suggesting potential architectural weaknesses. Moreover, AQSNet’s attempt at zero-shot transferring application clearly reveal its disability in tackling unseen segmentation masks (dashed rectangles in Fig. 2).
The limitations observed in existing DL-based unsupervised SQA methods—whether score prediction networks or pixel-level error localization networks are fundamentally rooted in common data characteristics and inherent conceptual or architectural deficiencies. From a data perspective, regardless of model generated or manual annotated, a large portion of erroneous pixels in unchecked segmentation masks are concentrated near object edges (see Section. IV-B3). Boundary neighbourhoods are often complex regions where precise classification into TP, FP, TN, or FN is inherently difficult. Secondly, the distribution of the four quality categories is typically highly imbalanced, often with vast background areas (TN), large correctly segmented foregrounds (TP), and much rarer instances of errors (FP, FN). This imbalance complicates the learning of discriminative features for all four classes, particularly the minority error classes, hindering a panoramic assessment. Thirdly, the spatial patterns and distributions of these four categories—the overall ’quality map’—can differ significantly between segmentation masks generated by humans and those produced by different automated models. Human-source segmentation masks are more likely to contain instance-wise deviation from ground truth than model-source ones as many of them are actually established based on historic inventory data. This variation in the underlying quality characteristics poses a significant challenge to the cross-domain transferability of any assessment model. Conceptually and architecturally, both prior paradigms exhibit flaws. Score prediction inherently lacks spatial detail. The theoretical basis of error localization methods, focusing solely on binary error detection or separate FP/FN localization, is also problematic for comprehensive assessment. A complete quality evaluation arguably requires considering all four pixel categories—true positive (TP), false positive (FP), true negative (TN), and false negative (FN)—as they constitute the full assessment result and provide mutual contextual priors that can aid identification. Additionally, the prevalent practice in prior work of simply concatenating image and mask features represents a naive interaction scheme (see Fig. 10). Compounded by the limited discriminative power of networks used previously, this often fails to effectively integrate information and prioritize assessment-relevant contexts, contributing to issues like class overlap observed in PEL methods. Finally, the reliance on traditional CNN-based backbones in many earlier methods limits their generalization capabilities, hindering zero-shot transferability to unseen scenes.
These fundamental challenges associated with SP and PEL networks suggest that significant advancements require moving beyond these approaches. This motivates the exploration of a third, distinct paradigm: Panoramic Quality Mapping (PQM), which aims for a comprehensive, pixel-wise classification across all quality categories (TP, FP, TN, and FN) to generate a complete quality map of the segmentation mask under evaluation. To realize this PQM vision and alleviate the interconnected issues of prior approaches, we propose SegAssess—a novel deep learning framework for pixel-level unsupervised segmentation quality assessment designed to achieve finer categorization, improved accuracy, and robust transferability. SegAssess embodies the PQM paradigm by explicitly formulating the SQA task as a four-class panoramic segmentation problem. It classifies every pixel within the segmentation mask under evaluation into one of four categories: true positive (TP, correctly identified foreground), false positive (FP, wrongly identified foreground), true negative (TN, correctly identified background), or false negative (FN, missed foreground), thus generating a complete quality map. Several key innovations differentiate SegAssess and enable effective PQM. To address the naive input fusion of previous methods, SegAssess establishes a promptable backbone based on the Segment Anything Model (SAM). This architecture uniquely treats the input segmentation mask as a prompt, integrating its embeddings with image features more effectively via cross-attention. Inheriting universal visual knowledge from the large-scale pre-trained SAM and employing remote sensing-specific fine-tuning further enhances domain adaptation and robust zero-shot capabilities. To specifically tackle challenges related to edge complexity and class imbalance identified earlier, SegAssess incorporates an Edge Guided Compaction (EGC) branch containing an Aggregated Semantic Filter (ASF) module. The EGC branch predicts object edges to focus computation on these complex regions, improving the localization of the rarer FP and FN classes while mitigating imbalance effects (see Fig 6); the ASF module enhances semantic filtering for better context utilization and edge prediction accuracy. Through RS-specific fine-tuning, the framework achieves domain adaptation while preserving SAM’s robust zero-shot capabilities. Finally, to directly combat poor cross-domain transferability arising from diverse mask sources, an Augmented Mixup Sampling (AMS) strategy is employed during training. AMS integrates multi-source segmentation masks, forcing the model to learn domain-agnostic features and addressing the dataset specificity that limits existing methods. Collectively, these innovations—grounded in the PQM paradigm—resolve the core deficiencies identified in prior SQA methods, namely coarse or incomplete categorization, spatial ambiguity in error localization, and poor transferability. To sum up, we provide contributions as follows:
@@
    We introduce SegAssess, a novel framework for pixel-level, unsupervised segmentation quality assessment based on the Panoramic Quality Mapping (PQM) paradigm. Compared to prior score prediction and pixel-level error localization methods, SegAssess offers finer assessment granularity (TP/FP/TN/FN), superior performance, and significantly enhanced transferability.
@@
    We design an Edge Guided Compaction (EGC) branch, incorporating a novel Aggregated Semantic Filter (ASF) module. These components specifically target complex edge regions, improving the model’s ability to differentiate challenging classes (particularly rare FP/FN pixels) more effectively than previous approaches.
@@
    We propose a dual strategy to achieve robust generalization and transferability. SegAssess leverages the domain-agnostic visual representations of a pre-trained SAM via an optimized prompt-based mechanism for strong zero-shot capabilities. This is complemented by an Augmented Mixup Sampling (AMS) strategy during training, which integrates multi-source segmentation data to explicitly enhance cross-domain robustness, overcoming a key limitation of prior work.
  • (4)
    We demonstrate through comprehensive experiments across diverse datasets that SegAssess achieves state-of-the-art (SOTA) performance under standard training-validation protocols. Crucially, it also exhibits remarkable zero-shot transferability to entirely unseen segmentation masks (both model- and human-sourced), establishing it as a new benchmark for unsupervised SQA.
@@ II-C2 Deep learning based methods
The advent of deep learning has spurred innovations in unsupervised SQA frameworks[34, 35, 36, 38, 39]. These typically involve training a model (often supervised by GT during development) to assess segmentation quality using the image and the segmentation mask as inputs, enabling reference-free assessment during deployment. Current methodologies fall into two categories, namely score prediction (SP) networks[34, 38, 39] and pixel-wise error localization (PEL) networks[35, 36]. The former paradigm aims at mapping input images and segmentation masks to quality scores. For instance,[38] rotated an input medical image with four angles and then design four parallel decoders to generate corresponding segmentation masks, which are utilized to collaboratively predict the final quality score through a regression model.[39] built the RS-SQA model that predicts the quality score of input segmentation masks based on a vision language model (VLM). While these SP models enable reference-free and automatic execution, they face two critical limitations. on the one hand, most of them adopt quantitative metrics that defined according to expert perception, thus there are still implicit uncertainty caused by the subjective design bias. On the other hand, score based assessment is implemented at image or object levels, restricting the analysis granularity. Alternatively, some studies pioneered the PEL paradigm that is introduced to address the granularity limitations of SP methods.[36] employed a U-net based network to process concatenated medical images and segmentation masks to identify wrong pixels. AQSNet[35] adopts a HRNet based architecture to process remote sensing images and segmentation masks to segment wrong and missing pixels. Though making processing, these PEL methods still suffer from significant drawbacks: they provide an incomplete assessment by typically ignoring correctly classified pixels (TP, TN), can exhibit spatial ambiguity (e.g., overlapping FP/FN predictions), and often demonstrate inadequate performance and poor transferability, especially to unseen mask sources.
@@ III-A Promptable HQ-SAM backbone
A core challenge in designing models for SQA is effectively integrating the image $I^{3\times H\times W}$ and the segmentation mask under evaluation $S^{H\times W}$ , where $H$ and $W$ are height and width. Prior methods often resort to simple channel-wise concatenation of the image and mask (or their features). This naive fusion approach, lacking specific mechanisms for multi-modal interaction, can hinder the model’s ability to treat the mask as a distinct subject for quality analysis rather than just another input channel, potentially leading to suboptimal feature interpretation.
To solve this problem, SegAssess employs a promptable backbone based on HQ-SAM. This architecture is adapted to treat the input segmentation mask $S^{H\times W}$ as a dense prompt, enabling sophisticated feature integration with the image representation $I^{3\times H\times W}$ via cross-attention mechanisms. As shown in Fig. 3, this backbone consists of three main components: an image encoder, a prompt encoder specifically tailored for mask inputs, and a high-quality decoder responsible for generating the PQM prediction.
Image encoder in SegAssess is similar to raw SAM that processes the input image $I^{3\times H\times W}$. It contains convolutional patchification operation and four attention stages. Each stage consists of a few Transformer based layers, which sequentially combines self-attention operation and multi-layer perceptron (MLP). Receiving an image $I^{3\times H\times W}$, stages of image encoder produces four corresponding intermediate embeddings ${F_{i}}^{{D_{im}}\times\frac{H}{16}\times\frac{W}{16}}$ where $i\in\{1,2,3,4\}$, $D_{im}$ means dimension number. To prepare for integrating with prompt features, the ${F_{4}}^{{D_{im}}\times\frac{H}{16}\times\frac{W}{16}}$ is compressed by two $3\times 3$ convolutions and obtain the high-level image embedding ${F_{im}}^{{D_{pr}}\times\frac{H}{16}\times\frac{W}{16}}$ where $D_{pr}$ denotes the channel number of prompt embeddings. The intermediate output from the first stage, ${F_{1}}^{{D_{im}}\times\frac{H}{16}\times\frac{W}{16}}$, is also retained for later use in the decoder.
@@
High-quality decoder is employed to incorporate tokenized images and segmentation masks to produce the final four-class PQM assessment map. Specifically, the image embedding ${F_{im}}^{{D_{pr}}\times\frac{H}{16}\times\frac{W}{16}}$ and prompt features ${F_{pr}}^{{D_{pr}}\times\frac{H}{16}\times\frac{W}{16}}$ are firstly added element-wise. Positional embeddings are incorporated, and the result is flattened to form the initial visual descriptors ${F}^{{D_{pr}}\times(\frac{H}{16}\times\frac{W}{16}})$. Additionally, to enhance the low-level perception, both the feature map ${F_{1}}^{{D_{im}}\times\frac{H}{16}\times\frac{W}{16}}$ from the first image encoder stage and the final image representation ${F_{im}}^{{D_{pr}}\times\frac{H}{16}\times\frac{W}{16}}$ are upsampled by two transposed convolution layers, and then sum up to obtain high-quality image representations ${F_{hq}}^{{D_{pr}}\times\frac{H}{4}\times\frac{W}{4}}$. Subsequently, Two sets of learnable output tokens are initialized: standard output tokens ${O}^{4\times{D_{pr}}}$ representing the four PQM classes (TP, FP, TN, FN), and high-quality output tokens ${O_{hq}}^{4\times{D_{pr}}}$ following the HQ-SAM [62] design to facilitate refined predictions. Mathematically, the final output tokens ${O_{f}}^{8\times{D_{pr}}}$ are yielded obeying $O_{f}=PE[O,O_{hq}]$ where $PE$ and $\cdot$] refer to the positional embedding and concatenation operation, respectively.
After obtaining ${F}^{{D_{pr}}\times\frac{H}{16}\times\frac{W}{16}}$ and ${O_{f}}^{8\times{D_{pr}}}$, features are propagated between them using two iterations of a Two Way Attention block. Each block consists of self-attention (SA) on the tokens, cross-attention (CA) from tokens to image features, an multi-layer perception (MLP), and another cross-attention from image features back to tokens (using the standard scaled dot-product attention mechanism, Eq.1):

### v2

@@ Abstract
High-quality image segmentation is fundamental to pixel-level geospatial analysis in remote sensing, necessitating robust segmentation quality assessment (SQA), particularly in unsupervised settings lacking ground truth. Although recent deep learning (DL) based unsupervised SQA methods show potential, they often suffer from coarse evaluation granularity, incomplete assessments, and poor transferability. To overcome these limitations, this paper introduces Panoramic Quality Mapping (PQM) as a new paradigm for comprehensive, pixel-wise SQA, and presents SegAssess, a novel deep learning framework realizing this approach. SegAssess distinctively formulates SQA as a fine-grained, four-class panoramic segmentation task, classifying pixels within a segmentation mask under evaluation into true positive (TP), false positive (FP), true negative (TN), and false negative (FN) categories, thereby generating a complete quality map. Leveraging an enhanced Segment Anything Model (SAM) architecture, SegAssess uniquely employs the input mask as a prompt for effective feature integration via cross-attention. Key innovations include an Edge Guided Compaction (EGC) branch with an Aggregated Semantic Filter (ASF) module to refine predictions near challenging object edges, and an Augmented Mixup Sampling (AMS) training strategy integrating multi-source masks to significantly boost cross-domain robustness and zero-shot transferability. Comprehensive experiments demonstrate that SegAssess achieves state-of-the-art (SOTA) performance and exhibits remarkable zero-shot transferability to unseen masks. The code is available at https://github.com/Yangbn97/SegAssess
@@ I Introduction
The proliferation of high-resolution remote sensing imagery has revolutionized geospatial analysis, enabling unprecedented spatial detail for Earth observation[1, 2]. Image segmentation (IS), which partitions pixels into semantically meaningful regions, serves as the foundational technique[3, 4, 5]. As the basis of pixel-level geospatial analysis, the segmentation quality remarkably impacts the reliability and robustness of the entire workflow[1, 2]. This critical dependency underscores the urgent demand for robust segmentation quality assessment (SQA) frameworks, especially in unsupervised settings where ground truth is unavailable.
@@
As shown in Fig. 1, existing SQA methods broadly fall into two categories: human-centric subjective evaluation and algorithm-based objective evaluation. Human-centric assessment[6, 7], while theoretically nuanced, is often economically impractical for large-scale applications. Supervised objective methods[8, 9, 10, 11, 12, 13] provide standardized evaluations by comparing outputs against ground truth (GT) references, but their fundamental reliance on GT data restricts their applicability in many real-world scenarios. Traditional unsupervised approaches[14, 15, 16] eliminate this dependency by analyzing intrinsic data patterns, but often struggle with transferability to new tasks or datasets.
With the emergence of deep learning, unsupervised SQA has advanced significantly, primarily following two paradigms: score prediction (SP)[17, 18, 19] and Pixel-level Error Localization (PEL)[20, 21]. SP networks process images and segmentation masks to predict a predefined quality score in a regression manner (Fig. 1). Though effective in certain contexts, these approaches assess quality at a coarse image or object level, lacking the pixel-wise diagnostic capabilities needed to localize specific defects.
@@
To address the granularity limitations of SP models, recent research has explored PEL networks[20, 21]. These methods attempt to obtain pixel-wise, objective assessment results by directly pinpointing per-pixel inaccuracies within the segmentation mask (Fig. 1). For example, one prior work[21] employed ”correct/incorrect” binary segmentation to simply identify erroneous pixels. Meanwhile, AQSNet[20] was designed to generate separate masks identifying false positive (FP, wrong foreground) and false negative (FN, missing foreground) pixels within the mask under evaluation. However, these PEL approaches, despite offering finer spatial detail, introduce their own set of limitations. Binary error classification (e.g., ”correct/incorrect”) lacks the necessary nuance to distinguish between different types of mistakes. Even methods like AQSNet that differentiate FP and FN errors face significant drawbacks. As illustrated by our preliminary experiments(Fig 2), AQSNet provides an incomplete assessment by overlooking correct categories, i.e.true positive (TP) and true negative (TN). Furthermore, its strategy of predicting FP and FN independently can lead to spatially overlapping error regions (pink areas in Fig. 2), suggesting potential architectural weaknesses. Moreover, AQSNet’s attempt at zero-shot transferring application clearly reveals its disability in tackling unseen segmentation masks, particularly when the error characteristics differ from the training data (dashed rectangles in Fig. 2).
The limitations observed in existing DL-based SQA methods are fundamentally rooted in common data characteristics and inherent conceptual or architectural deficiencies. From a data perspective, erroneous pixels are often concentrated near complex object edges, and the four quality categories (TP, FP, TN, FN) are typically highly imbalanced, hindering a panoramic assessment (see Section IV-B3). Architecturally, prior paradigms exhibit flaws. Score prediction inherently lacks spatial detail, while the theoretical basis of error localization methods is problematic for comprehensive assessment. A complete quality evaluation arguably requires considering all four pixel categories, as they provide mutual contextual priors that can aid identification (see Section V-B).
These fundamental challenges motivate the exploration of a third, distinct paradigm: Panoramic Quality Mapping (PQM). We define PQM as the reformulation of unsupervised SQA into a comprehensive, pixel-wise classification task, wherein every pixel of the mask under evaluation is classified into one of four fundamental quality categories—true positive (TP), false positive (FP), true negative (TN), or false negative (FN)—thereby generating a complete quality map. This paradigm aims for a holistic assessment that resolves the issues of incomplete categorization and spatial ambiguity found in prior methods.
To realize this PQM vision and alleviate the interconnected issues of prior approaches, we propose SegAssess—a novel deep learning framework for pixel-level unsupervised segmentation quality assessment designed to achieve finer categorization, improved accuracy, and robust transferability. SegAssess embodies the PQM paradigm by explicitly formulating the SQA task as a four-class panoramic segmentation problem. It is built upon an High-quality Segment Anything Model[22] (HQ-SAM) architecture, uniquely treating the input segmentation mask as a prompt to facilitate effective feature integration via cross-attention. Key innovations include an Edge Guided Compaction (EGC) branch containing an Aggregated Semantic Filter (ASF) module to specifically tackle challenges related to edge complexity and class imbalance. Furthermore, to directly combat poor cross-domain transferability, an Augmented Mixup Sampling (AMS) strategy is employed during training.
To sum up, we provide contributions as follows:
@@
    We introduce SegAssess, a novel framework for pixel-level, unsupervised segmentation quality assessment based on the Panoramic Quality Mapping (PQM) paradigm. Compared to prior methods, SegAssess offers finer assessment granularity (TP/FP/TN/FN), superior performance, and significantly enhanced transferability.
@@
    We design an Edge Guided Compaction (EGC) branch, incorporating a novel Aggregated Semantic Filter (ASF) module. These components specifically target complex edge regions, improving the model’s ability to differentiate challenging classes more effectively.
@@
    We propose a dual strategy to achieve robust generalization and transferability. SegAssess leverages the domain-agnostic visual representations of a pre-trained SAM via an optimized prompt-based mechanism and is complemented by an Augmented Mixup Sampling (AMS) strategy during training.
@@ II-C2 Deep learning based methods
The advent of deep learning has spurred innovations in unsupervised SQA frameworks[17, 20, 21, 18, 19]. These typically involve training a model (often supervised by GT during development) to assess segmentation quality using the image and the segmentation mask as inputs, enabling reference-free assessment during deployment. Current methodologies fall into two categories, namely score prediction (SP) networks[17, 18, 19] and pixel-wise error localization (PEL) networks[20, 21]. The former paradigm aims at mapping input images and segmentation masks to quality scores. For instance, the method proposed in [18] rotated an input medical image with four angles and then design four parallel decoders to generate corresponding segmentation masks, which are utilized to collaboratively predict the final quality score through a regression model. RS-SQA model[19] predicts the quality score of input segmentation masks based on a vision language model (VLM). While these SP models enable reference-free and automatic execution, they face two critical limitations. on the one hand, most of them adopt quantitative metrics that defined according to expert perception, thus there are still implicit uncertainty caused by the subjective design bias. On the other hand, score based assessment is implemented at image or object levels, restricting the analysis granularity. Alternatively, some studies pioneered the PEL paradigm that is introduced to address the granularity limitations of SP methods. A previous work[21] employed a U-net based network to process concatenated medical images and segmentation masks to identify wrong pixels. AQSNet[20] adopts a HRNet based architecture to process remote sensing images and segmentation masks to segment wrong and missing pixels. Though making processing, these PEL methods still suffer from significant drawbacks: they provide an incomplete assessment by typically ignoring correctly classified pixels (TP, TN), can exhibit spatial ambiguity (e.g., overlapping FP/FN predictions), and often demonstrate inadequate performance and poor transferability, especially to unseen mask sources.
@@ III-A Promptable HQ-SAM backbone
A core challenge in designing models for SQA is effectively integrating the image $I^{3\times H\times W}$ and the segmentation mask under evaluation $S^{H\times W}$ , where $H$ and $W$ are height and width. Prior methods often resort to simple channel-wise concatenation, which can hinder feature interpretation. To solve this problem, SegAssess employs a promptable backbone adapted from HQ-SAM[22]. Our key adaptation is not a re-design of the backbone’s internal layers, but its novel application to the SQA task, where we uniquely treat the input mask as a dense prompt, enabling sophisticated feature integration with the image representation via cross-attention mechanisms. This approach is motivated by the powerful multi-modal fusion capabilities inherent in large foundation models like SAM. By leveraging the model’s pre-trained weights, we treat the image and the mask as two distinct modalities and capitalize on the backbone’s ability to fuse them effectively via cross-attention. This strategy has been shown to be more robust and effective than adapting a smaller, single-modality network to multi-model domain, allowing for a more nuanced integration of information for the final assessment. The backbone consists of an image encoder, a prompt encoder, and a high-quality decoder.
Image encoder follows the standard Vision Transformer (ViT) architecture used in SAM to process the input image $I^{3\times H\times W}$ through several attention stages and consequently produces multi-scale feature embeddings ${F_{i}}^{{D_{im}}\times\frac{H}{16}\times\frac{W}{16}}$ where $i\in\{1,2,3,4\}$, $D_{im}$ means dimension number. The final high-level image embedding from the last stage ${F_{4}}^{{D_{im}}\times\frac{H}{16}\times\frac{W}{16}}$ is dimensionally compressed to obtain ${F_{im}}^{{D_{pr}}\times\frac{H}{16}\times\frac{W}{16}}$, which matches the prompt embedding channel number $D_{pr}$. Concurrently, the first-stage embedding ${F_{1}}^{{D_{im}}\times\frac{H}{16}\times\frac{W}{16}}$ is retained to provide essential low-level details to the decoder.
@@
High-quality decoder is employed to fuse tokenized images and segmentation masks to produce the final four-class PQM assessment map.
Specifically, the image embedding ${F_{im}}^{{D_{pr}}\times\frac{H}{16}\times\frac{W}{16}}$ and prompt features ${F_{pr}}^{{D_{pr}}\times\frac{H}{16}\times\frac{W}{16}}$ are firstly added element-wise. Positional embeddings are incorporated, and the result is then flattened and transposed to form the initial visual descriptors ${F}^{(\frac{H}{16}\times\frac{W}{16})\times{D_{pr}}}$. Additionally, to enhance the low-level perception, both the feature map ${F_{1}}^{{D_{im}}\times\frac{H}{16}\times\frac{W}{16}}$ from the first image encoder stage and the final image representation ${F_{im}}^{{D_{pr}}\times\frac{H}{16}\times\frac{W}{16}}$ are upsampled by two transposed convolution layers, and then sum up to obtain high-quality image representations ${F_{hq}}^{{D_{pr}}\times\frac{H}{4}\times\frac{W}{4}}$.
Subsequently, Two sets of learnable output tokens are initialized: standard output tokens ${O}^{4\times{D_{pr}}}$ representing the four PQM classes (TP, FP, TN, FN), and high-quality output tokens ${O_{hq}}^{4\times{D_{pr}}}$ following the HQ-SAM [22] design to facilitate refined predictions. Mathematically, the final output tokens ${O_{f}}^{8\times{D_{pr}}}$ are yielded obeying $O_{f}=PE[O,O_{hq}]$ where $PE$ and $\cdot$] refer to the positional embedding and concatenation operation, respectively.
After obtaining ${F}^{(\frac{H}{16}\times\frac{W}{16})\times{D_{pr}}}$ and ${O_{f}}^{8\times{D_{pr}}}$, features are propagated between them using two iterations of a Two Way Attention block. Each block consists of self-attention (SA) on the tokens, cross-attention (CA) from tokens to image features, an multi-layer perception (MLP), and another cross-attention from image features back to tokens (using the standard scaled dot-product attention mechanism, Eq.1):